OpenStudy (anonymous):

Basic calc, but I can't seem to figure it out! f'(x)∈ℝ,f(1)=1,∫xf'(x)dx from 0 to 1 is 1. Find ∫f(x)dx from 0 to 1. I'd like to see the step-by-step method, please.

OpenStudy (anonymous):

Have you tried using integration by parts?

OpenStudy (anonymous):

∫xf'(x)dx=xf(x)-∫f'(x)dx

OpenStudy (anonymous):

Facepalm moment, haha. Thanks.

OpenStudy (anonymous):

.-)

OpenStudy (turingtest):

actually I think andras made a little mistake...\[\int_{0}^{1} xf'(x)=xf(x)|_{0}^{1}-\int_{0}^{1} f(x)dx=1\]\[\int_{0}^{1} f(x)dx=xf(x)|_{0}^{1}-1=f(1)-1=1-1=0\]

OpenStudy (phi):

In case it isn't obvious from the posted answers, this question is about integration by parts \[\int\limits_{}^{}u \text{ }dv=uv-\int\limits_{}^{}v \text{ }du\] Here, let x= u and f'(x) dx = dv

OpenStudy (anonymous):

Yeah, all I needed was for him to tell me integrate by parts. Honestly, I didn't read his actual response except that. Heh. The steps > the answer.

OpenStudy (anonymous):

Also, I can't believe I forgot about this. I took calc just two semesters ago. >.>